# Universal bound on the Lyapunov spectrum of quantum master equations

### v1

@@ 4.2 Condition on the embedded classical master equation imposed by the Schwarz-class
It can be shown (see, e.g., the discussion in [ChKiMu2024] and references therein) that the convex set of positive maps on $\mathcal{M}_{d}(\mathbb{C})$ yields a hierarchy of inclusions finer than (12). For the present work we only notice that between the convex sets of positive and $2$-positive maps sits the set $\mathcal{S}_{1}$ (also convex) of Schwarz maps:
$\mathcal{P}\,\equiv\,\mathcal{P}_{\mathfrak{1}}\supset\mathcal{S}_{1}\supset\mathcal{P}_{\mathfrak{2}}$
We characterize Schwarz maps by considering their adjoint with respect to the Hilbert-Schmidt inner product. Specifically, if $\bm{X}$ is an arbitrary linear map, we define its Hilbert-Schmidt adjoint $\bm{X}^{\ddagger}$ by requiring
$\operatorname{Tr}\left(\operatorname{A}^{\dagger}\bm{X}[\operatorname{B}]\right)=\operatorname{Tr}\left(\big(\bm{X}^{\ddagger}[\operatorname{A}]\big)^{\dagger}\operatorname{B}\right)\qquad\forall\,\operatorname{A},\operatorname{B}\in\mathcal{M}_{d}(\mathbb{C})$
@@
$\bm{X}^{\ddagger}[\operatorname{1}_{d}]=\operatorname{1}_{d}\,.$
@@
$\bm{X}^{\ddagger}[\operatorname{A}^{\dagger}\operatorname{A}]\,\geq\,\bm{X}^{\ddagger}[\operatorname{A}^{\dagger}]\bm{X}^{\ddagger}[\operatorname{A}]\qquad\forall\,\operatorname{A}\,\in\,\mathcal{M}_{d}(\mathbb{C})$
@@
Let us now suppose that that (5) generates a semigroup in $\mathcal{S}_{1}$. Then we claim that the classical master equation constructed out of any of the particular solutions is such that
@@ 4.2.1 Proof of the claim
$\operatorname{P}_{i,j}(t)=\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\bm{\phi}^{\scriptscriptstyle{(j)}}{}^{\dagger}(t)$
@@
$\mathfrak{L}(t)^{\ddagger}[\operatorname{P}_{j,j}(t)]\,\geq\,\operatorname{P}_{i,j}^{\dagger}(t)\mathfrak{L}^{\ddagger}(t)[\operatorname{P}_{i,j}(t)]+\mathfrak{L}(t)^{\ddagger}[\operatorname{P}_{i,j}^{\dagger}(t)]\operatorname{P}_{i,j}(t)\,.$
From the left-hand side of the inequality we get
$\sum_{i=1}^{d}\sum_{j\neq i}^{d}\operatorname{Tr}\left(\operatorname{P}_{j,j}(t)\mathfrak{L}(t)^{\ddagger}\big[\operatorname{P}_{j,j}(t)\big]\right)=\sum_{i,j=1}^{d}(1-\delta_{i,j})\operatorname{Tr}\left(\mathfrak{L}(t)\big[\operatorname{P}_{j,j}(t)\big]\operatorname{P}_{j,j}(t)\right)=(d-1)\,\operatorname{Tr}\mathsf{W}(t)$
@@
$\sum_{i=1}^{d}\sum_{j\neq i}^{d}\operatorname{Tr}\left(\operatorname{P}_{j,j}(t)\big(\operatorname{P}_{i,j}^{\dagger}(t)\mathfrak{L}^{\ddagger}(t)[\operatorname{P}_{i,j}(t)]+\mathfrak{L}(t)^{\ddagger}[\operatorname{P}_{i,j}^{\dagger}(t)]\operatorname{P}_{i,j}(t)\big)\right)$
@@
$=\sum_{i,j=1}^{d}(1-\delta_{i,j})\operatorname{Tr}\left(\operatorname{P}_{j,i}(t)\mathfrak{L}^{\ddagger}(t)[\operatorname{P}_{i,j}(t)]+\mathfrak{L}(t)^{\ddagger}[\operatorname{P}_{i,j}^{\dagger}(t)]\operatorname{P}_{i,j}(t)\right)=2\,\left(\operatorname{Tr}\mathscr{L}(t)-\operatorname{Tr}\mathsf{W}(t)\right)$
The claim follows at once by comparing the two above identities.
@@ 5 Main result
The Lyapunov characteristic exponents $\big\{\lambda_{i}\big\}_{i=0}^{d^{2}-1}$ (see, e.g., [OseV1968, BeGaGiSt1980] or the monographs [CeCeVu2009, PiPo2016]; a brief summary is also in  LABEL:app:Lyapunov below) of a quantum master equation specify the asymptotic properties of averages over a state operator. As such they are also experimentally accessible. The Lyapunov exponents of a trace- and self-adjointness preserving, contractive (also referred to as positive) semigroup solving (5) satisfy
$0\,\geq\,\lambda_{\mathfrak{0}}\,\geq\,\lambda_{\mathfrak{1}}\,\geq\,\dots\,\geq\,\lambda_{d^{2}-1}$
We are now in a position to prove results equivalent to those of [ChvEnKiMG2025, vEnChKiMG2025].
@@
If a semigroup generated by (5) is trace-preserving, self-adjointness preserving, contractive,
@@
    and also 22-positive, i.e., (23) holds, then the smallest Lyapunov exponent obeys the bound
@@
    and of Schwarz class, i.e., (25) holds, then
@@
    but does not enjoy any additional property, then the lower bound on the smallest Lyapunov exponent is trivial:
@@
Before turning to the proof some observations are in order.
  • The set of Lyapunov exponents identifies the asymptotic decay rates of a contractive semigroup via
    $\Gamma_{\mathfrak{i}}=-\lambda_{\mathfrak{i}}$ $\mathfrak{i}=0,\dots,d^{2}-1$
    In terms of the asymptotic decay rates the inequalities read
    $\Gamma_{d^{2}-1}\leq c_{d}\sum_{\ell=1}^{d^{2}-1}\Gamma_{\ell}$
    where the universal prefactor $c_{d}$ is
    $c_{d}=\begin{cases}\end{cases}$

### v2

@@ 4.2 Condition on the embedded classical master equation imposed by the Schwarz-class
It can be shown (see, e.g., the discussion in [ChKiMu2024] and references therein) that the convex set of positive maps on $\mathcal{M}_{d}(\mathbb{C})$ yields a hierarchy of inclusions finer than (23). For the purposes of the present work, we only note that between the convex sets of positive and $2$-positive maps lies the set $\mathcal{S}_{1}$ (also convex) of Schwarz maps:
$\mathcal{P}\,\equiv\,\mathcal{P}_{1}\supset\mathcal{S}_{1}\supset\mathcal{P}_{2}$ (53)
We characterize Schwarz maps by considering their adjoint with respect to the Hilbert-Schmidt inner product (37). For an arbitrary linear map $\bm{X}$, we define its Hilbert-Schmidt adjoint $\bm{X}^{\ddagger}$ by requiring
$\langle\operatorname{A}\,,\bm{X}[\operatorname{B}]\rangle:=\operatorname{Tr}\left(\operatorname{A}^{\dagger}\bm{X}[\operatorname{B}]\right)=\operatorname{Tr}\left(\big(\bm{X}^{\ddagger}[\operatorname{A}]\big)^{\dagger}\operatorname{B}\right)=:\langle\bm{X}^{\ddagger}[\operatorname{A}]\,,\operatorname{B}\rangle\qquad\forall\,\operatorname{A},\operatorname{B}\in\mathcal{M}_{d}(\mathbb{C})$ (54)
@@
$\bm{X}^{\ddagger}[\operatorname{1}_{d}]=\operatorname{1}_{d}\,.$ (55)
@@
$\bm{X}^{\ddagger}[\operatorname{A}^{\dagger}\operatorname{A}]\,\geq\,\bm{X}^{\ddagger}[\operatorname{A}^{\dagger}]\bm{X}^{\ddagger}[\operatorname{A}]\qquad\forall\,\operatorname{A}\,\in\,\mathcal{M}_{d}(\mathbb{C})$ (56)
@@
Now, let us suppose that (7) generates a semigroup in $\mathcal{S}_{1}$. Then we claim that the classical master equation constructed from any of the particular solutions is such that
@@ 4.2.1 Proof of the claim
$\operatorname{P}_{i,j}(t)=\bm{\phi}_{i}(t)\bm{\phi}_{j}^{\dagger}(t)$ (59)
@@
$\mathfrak{L}^{\ddagger}(t)[\operatorname{P}_{j,j}(t)]\,\geq\,\operatorname{P}_{i,j}^{\dagger}(t)\mathfrak{L}^{\ddagger}(t)[\operatorname{P}_{i,j}(t)]+\mathfrak{L}^{\ddagger}(t)[\operatorname{P}_{i,j}^{\dagger}(t)]\operatorname{P}_{i,j}(t)\,.$ (60)
From the left-hand side of the inequality we obtain:
$\sum_{i=1}^{d}\sum_{j\neq i}^{d}\operatorname{Tr}\left(\operatorname{P}_{j,j}(t)\mathfrak{L}^{\ddagger}(t)\big[\operatorname{P}_{j,j}(t)\big]\right)=\sum_{i,j=1}^{d}(1-\delta_{i,j})\operatorname{Tr}\left(\mathfrak{L}(t)\big[\operatorname{P}_{j,j}(t)\big]\operatorname{P}_{j,j}(t)\right)=(d-1)\,\operatorname{Tr}\mathsf{W}(t)$ (61)
@@
$\sum_{i=1}^{d}\sum_{j\neq i}^{d}\operatorname{Tr}\left(\operatorname{P}_{j,j}(t)\big(\operatorname{P}_{i,j}^{\dagger}(t)\mathfrak{L}^{\ddagger}(t)[\operatorname{P}_{i,j}(t)]+\mathfrak{L}^{\ddagger}(t)[\operatorname{P}_{i,j}^{\dagger}(t)]\operatorname{P}_{i,j}(t)\big)\right)$
@@
$=\sum_{i,j=1}^{d}(1-\delta_{i,j})\operatorname{Tr}\left(\operatorname{P}_{j,i}(t)\mathfrak{L}^{\ddagger}(t)[\operatorname{P}_{i,j}(t)]+\mathfrak{L}^{\ddagger}(t)[\operatorname{P}_{i,j}^{\dagger}(t)]\operatorname{P}_{i,j}(t)\right)=2\,\left(\operatorname{Tr}\mathscr{L}(t)-\operatorname{Tr}\mathsf{W}(t)\right)$ (62)
The claim follows at once by comparing the two identities above.
5 A distilled view on Lyapunov exponents
We briefly recall the definition and the properties of Lyapunov exponents propaedeutic to the proof of our main result. In doing so, we mainly draw from [AdrL1995] and the more recent [BarL2017].
Let $\mathsf{X}(t)$ be the Cauchy matrix solution of a time-non-autonomous differential system on a $d$-dimensional Euclidean vector space $\mathbb{E}$ with the initial condition given by the identity matrix on $\mathbb{E}$:
$\begin{split}&\dot{\mathsf{X}}(t)=\mathsf{A}(t)\mathsf{X}(t)\\ &\mathsf{X}(0)=\mathsf{1}_{d}\end{split}$ (63)
We assume that $\mathsf{A}(t)$ is a continuous and uniformly bounded square matrix for all $t\,\geq\,0$. Under these hypotheses, if $\big\{\bm{g}_{i}\big\}_{i=1}^{d}$ is any basis of $\mathbb{E}$, then the characteristic exponents
$\chi(\bm{g}_{i})=\limsup_{t\to\infty}\frac{1}{t}\ln\|\mathsf{X}(t)\bm{g}_{i}\|_{a}$ $i=1,\dots,d$ (64)
are well defined quantities. The definition is independent of the choice of the norm $\left\|\cdot\right\|_{a}$ provided the norm is equivalent to the Euclidean norm which corresponds to $a$ equal to $2$. We elaborate some more on norm equivalence in Appendix LABEL:app:norms.
Lyapunov introduced the term “normal basis” for a basis $\big\{\bm{f}_{i}\big\}_{i=1}^{d}$ that minimizes the sum over all the characteristic exponents [OseV1968, BeGaGiSt1980]:
$\sum_{i=1}^{d}\chi(\bm{f}_{i})=\liminf_{t\to\infty}\frac{1}{t}\ln\det\mathsf{X}(t)=\liminf_{t\to\infty}\frac{1}{t}\int_{0}^{t}\mathrm{d}s\operatorname{Tr}\mathsf{A}(s)$ (65)
Correspondingly, the exponents (66) computed on a normal basis are called Lyapunov exponents:
$\lambda_{i}=\chi(\bm{f}_{i})$ $i=1,\dots,d$ (66)
Geometrically, a normal basis filters $\mathbb{E}$ into a flag of $n\,\leq\,d$ nested linear subspaces
$\{0\}:=\mathbb{V}_{0}\subset\mathbb{V}_{1}\subset\dots\subset\mathbb{V}_{n}=\mathbb{E}$ (67)
each of which is spanned by a subset of the elements of the basis. This means that each $\mathbb{V}_{i}$ corresponds to a distinct value of the Lyapunov exponents. Clearly, the flag has $d$ elements if the Lyapunov exponents are non-degenerate. In what follows, we label the exponents from one to $d$ to simplify the notation.
The QR decomposition of $\mathsf{X}(t)$, see e.g. [HoJo2013, Th 2.1.14], gives a more computationally concrete meaning to Lyapunov exponents. Namely, it is always possible to uniquely factorize the solution $\mathsf{X}(t)$ of (63) into a unitary matrix $\mathsf{U}(t)$ and an upper triangular matrix $\mathsf{T}(t)$ having positive diagonal entries:
$\mathsf{X}(t)=\mathsf{U}(t)\,\mathsf{T}(t)$ (68)
Upon denoting by $\mathsf{T}_{i,i}(t)$ the diagonal entries of $\mathsf{T}(t)$, we arrive at
$\lambda_{i}=\limsup_{t\to\infty}\frac{1}{t}\ln\mathsf{T}_{i,i}(t)$ $\forall\,i=1,\dots,d$ (69)
[…]
$\begin{split}&\dot{\mathsf{Y}}(t)=-\mathsf{A}^{\dagger}(t)\mathsf{Y}(t)\\ &\mathsf{Y}(0)=\mathsf{1}_{d}\end{split}$ (70)
Traditionally see e.g. [ByViGrNe1966, § 1.4] or [AdrL1995, § I.1], (70) is referred to as the “adjoint equation” to (63). To avoid confusion with the concept of “adjoint” used in section 4.2, we refer to (70) as the Lyapunov “dual system”. The Cauchy matrix solving (70) is the inverse of the adjoint of that solving (63).
[…]
The important consequence is the preservation of the inner product
$\left\langle\,\mathsf{Y}(t)\bm{w}\,,\mathsf{X}(t)\bm{v}\,\right\rangle=\left\langle\,\bm{w}\,,\bm{v}\,\right\rangle$ $\forall\bm{v}\,,\bm{w}\in\mathbb{E}_{d}$ (72)
Upon applying the Cauchy-Schwarz inequality we get
$\|\mathsf{Y}(t)\bm{w}\|_{2}^{2}\,\|\mathsf{X}(t)\bm{v}\|_{2}^{2}\geq\left|\left\langle\,\bm{w}\,,\bm{v}\,\right\rangle\right|^{2}$ $\forall\bm{v}\,,\bm{w}\in\mathbb{E}_{d}$ (73)
Hence, if we arrange the Lyapunov exponents of the primal dynamics (63) in decreasing order and those of the dual dynamics (70) in increasing order, and we choose $\bm{v}$ and $\bm{w}$ as the corresponding elements of normal bases spanning the two flags of nested subspaces, we generally obtain the inequalities:
$\lambda_{i}+\Gamma_{i}\,\geq\,0$ $i=1,\dots\,d$ (74)
These inequalities raise the question of identifying the conditions under which
$\lambda_{i}=-\Gamma_{i}$ $i=1,\dots,d$ (75)
Cardinal results in the theory of Lyapunov exponents due to Oseledets [OseV1968, OseV2008] and Ragunathan [RagM1979], see also [BeGaGiSt1980], show that answering this question is equivalent to determining the conditions under which we can replace $\limsup$ with $\lim$ in the definition (66), or equivalently (69), of the exponents. The answer identifies the class of Lyapunov “regular” systems, which enjoy the property that the limit
$\lim_{t\to\infty}\Big(\mathsf{X}^{\dagger}(t)\mathsf{X}(t)\Big)^{\frac{1}{2\,t}}=\mathsf{Q}$ (76)
exists and specifies a well-defined, self-adjoint, positive square matrix $\mathsf{Q}$. We refer to [BarL2017, Th 7.2.2] for a proof of the equivalence of the characterization of regularity (76) to the one used in [OseV1968, BeGaGiSt1980] and to [BarL2017, Th 4.3.1] for a proof that Lyapunov regularity is equivalent to (75).
When the limit (76) exists, the Lyapunov exponents coincide with the logarithms of the eigenvalues $\big\{\mathsf{q}_{i}\big\}_{i=1}^{d}$ of $\mathsf{Q}$:
$\lambda_{i}=-\Gamma_{i}=\ln\mathsf{q}_{i}$ $i=1,\dots,d$ (77)
Furthermore, for a regular system, the sum of the Lyapunov exponents exactly specifies the growth rate of a volume form:
$\sum_{i=1}^{d}\lambda_{i}=\lim_{t\to\infty}\frac{1}{t}\int_{0}^{t}\mathrm{d}s\operatorname{Tr}\mathsf{A}(s)$ (78)
The characterization of regularity (76) is the basis of many algorithms used in the physics literature [CeCeVu2009, PiPo2016] to compute Lyapunov exponents. Mathematically, (76) is a stronger requirement than global existence and uniqueness of solutions (63). Fortunately, it is possible to prove sufficient conditions for regularity only based on the properties of the matrix $\mathsf{A}(t)$. Specifically, (63) is Lyapunov regular if $\mathsf{A}(t)$, in addition to be continuous and bounded, enjoys one of the following properties:
  1. C-i
    is time independent
    $\mathsf{A}(t)=\mathsf{A}(0)=\mathsf{A}$ $\forall\,t\,\geq\,0$ (79)
    as we can directly verify by diagonalization see e.g. [AdrL1995, Lemma 3.5.3]);
  2. C-$i$
    is periodic in time with period $T>0$:
    $\mathsf{A}(t+T)=\mathsf{A}(t)$ $\forall t\,\geq\,0$ (80)
    This claim can be verified by means of Floquet theorem see e.g. [AdrL1995, Lemma 3.5.3].
  3. C-iii
    $\mathsf{A}(t)$ is almost reducible to a time independent or periodic matrix. Almost reducible means that for any $\delta\,\geq\,0$ there exists a nonsingular and continuously differentiable matrix $\mathsf{L}_{\delta}(t)$ such that
    $\mathsf{B}(t)+\mathsf{C}(t)=\mathsf{L}_{\delta}^{-1}(t)\mathsf{A}(t)\mathsf{L}_{\delta}(t)-\mathsf{L}_{\delta}^{-1}(t)\mathsf{\dot{L}}_{\delta}(t)$ (81)
    such that $\mathsf{B}(t)$ is either constant or periodic and
    $\|\mathsf{C}(t)\|\,\leq\,\delta$ (82)
    The matrix norm may be any norm induced by a vector norm equivalent to the Euclidean norm. The proof of this last sufficient condition is more laborious [AdrL1995, Th 3.5.2].
The sufficient conditions C-iC-iii encompass time-autonomous, periodic and almost periodic generators of quantum master equations. Generators typically, if not always, encountered in applications involving deterministic quantum master equations satisfy one of the sufficient conditions above.
More generally, Oseledets’ multiplicative ergodic theorem [OseV1968, RagM1979, BeGaGiSt1980, OseV2008] guarantees the existence of the limit (76) when, roughly speaking, the family of Cauchy matrices constitutes a statistically stationary, ergodic sequence of matrices. The theorem is of central importance for non-autonomous linear systems obtained by linearization around a solution $\bm{x}(t)$ of a non-linear differential equation to study its stability versus infinitesimal change of the initial data. In this latter case the hypotheses of Oseledets’ theorem are satisfied e.g. when $\bm{x}(t)$ is ergodic. Physics applications of Oseledets’ theorem to non-linear dynamics are beyond the scopes of the present contribution and are extensively discussed in [CeCeVu2009, PiPo2016]. We expect, however, Oseledets’ theorem to be relevant for studying ensembles of random quantum master equations [DeLaTaChZy2019].
@@ 6 Main result
We are now in a position to prove results equivalent to those of [ChvEnKiMG2025, vEnChKiMG2025]. Semigroups generated by (7) are always trace and self-adjointness preserving. In addition
@@
If the elements of the semigroup are all contractive and Lyapunov regular, the Lyapunov exponents satisfy
$0\,=\,\lambda_{0}\,\geq\,\lambda_{1}\,\geq\,\dots\,\geq\,\lambda_{d^{2}-1}$ (83)
while, according to the definition of section 5, the exponents of the dual system are
$\Gamma_{d^{2}-1}=-\lambda_{d^{2}-1}\,\geq\,\dots\,\geq\,\Gamma_{1}=-\lambda_{1}\,\geq\,\Gamma_{0}=-\lambda_{0}=0$ (84)
Furthermore,
@@
    if the semigroup is also $2$-positive, i.e., (44) holds, then the smallest Lyapunov exponent obeys the bound
@@
    if instead the semigroup is Schwarz class, i.e., (58) holds, then
@@
    finally, if the semigroup does not enjoy any additional property, then the lower bound on the smallest Lyapunov exponent is trivial:
@@
Before turning to the proof, some observations are in order.
  • Our hypotheses require contractivity for all the elements of the semigroup or, in other words, positive divisibility of the linear maps.
  • The dual exponents $\Gamma_{i}$’s specify the asymptotic decay rates of a contractive semigroup. The inequalities for the asymptotic decay rates read:
    $\Gamma_{d^{2}-1}\leq\varkappa_{d}\sum_{\ell=1}^{d^{2}-1}\Gamma_{\ell}$ (88)
    where the universal prefactor $\varkappa_{d}$ is
    $\varkappa_{d}=\begin{cases}\end{cases}$